# Fast Decentralized State Estimation for Legged Robot Locomotion via EKF and MHE

### v1

@@ I Introduction
Due to the inherently underactuated nature of legged robots, an accurate and high-frequency estimation of the torso velocity and orientation [2] is essential for the control of dynamic locomotion. This is a ubiquitous requirement for all dynamic legged robots regardless of how many legs they have. Extended Kalman filter (EKF) and its variants [2, 3, 4] are commonly used to solve this problem. These methods provide an accurate recursive solution for the fusion of onboard proprioceptive sensors including Inertial Measurement Units (IMU) and joint encoders. However, when working with locomotion behaviors associated with longer aerial phases, measurements from leg kinematics cannot be used to correct the floating base states. With only inertial sensors, the estimation of velocities can quickly diverge. Additionally, the yaw drift is inevitable when only using proprioceptive sensors [2]. To address this problem, several works [5, 6] have included exteroceptive sensors such as cameras and Lidar within the filtering methods. However, [5] primarily focused on pose drift correction, and [6] integrated vision and IMU into a single measurement, which compromised the robustness of the estimation due to potential correlations.
@@
To address this limitation, smoothing method that utilizes windowed optimization to include a history of sensor measurements is advantageous since multirate sensor data can be fused in the window. [7] utilizes a factor graph method, which tightly couples vision with IMU and encoder measurements through preintegration. However, with huge computation demands, its estimation rate remains insufficient for real-time control. Another smoothing method, Moving Horizon Estimation (MHE), known as the dual problem of Model Predictive Control (MPC), provides an alternative formulation of the windowed optimization. Similar to MPC [8], when the dynamics is nonlinear, the MHE becomes a nonconvex optimization which yields demanding computation. Inspired by MHE, [9] proposed a Quadratic Program (QP) based state estimator using full-body dynamics with state constraints, to estimate contact force and joint torques. An inequality-constrained MHE was explicitly applied in [10] for a bipedal robot, demonstrating its ability to handle constraints and non-Gaussian noises. However, only proprioceptive sensors were considered in [10], and old measurements prior to the smoothing window were directly discarded, which compromised the optimality of the estimation.
@@
Figure 1: The proposed estimation framework and the application on the dynamic hopping robot PogoX.
@@
In this paper, we aim to develop a MHE based framework that achieves multirate sensor fusion of both proprioceptive and exteroceptive sensors, providing fast and accurate estimation for legged robot control. The computation burden of nonlinear optimization of the canonical MHE is addressed by the decentralization of the orientation and velocity estimation: the nonlinear orientation estimation replies on an EKF, whereas the linear velocity estimation is solved via the MHE. The framework is applicable to any legged systems that have an IMU, leg joint encoders, and cameras. The work has the following contributions:
  • A formulation of a novel state estimation framework using MHE, which fuses multirate proprioceptive and exteroceptive sensors and incorporates state constraints.
  • A decentralization of the legged state estimation to a lightweight orientation EKF and a linear MHE for linear velocity estimation, achieving real-time performance.
@@
  • Evaluations of the proposed decentralized estimators on various legged robotics systems, including the unipedal hopping robot PogoX [12], bipedal robot Cassie [3], and quadrupedal robot Unitree Go1 [13].
@@ II Related Work
State-of-the-art methods for legged robot state estimation are categorized into filtering and smoothing approaches. For filtering approaches, they typically perform fusions of proprioceptive sensors (IMUs, force/torque sensors, and joint encoders) with high frequency (100 - 2000 Hz) using Kalman Filters (KFs). [2] showed that using EKF, the robot pose, velocity, and IMU bias could be estimated on quadrupedal robots. Invariant EKF [3] was proposed to improve the convergence of the estimation. The filtering approaches have also been widely implemented on bipedal robots [4]. While providing high-frequency velocity estimation, the estimation accuracy is strongly affected by the availability of foot contact and accurate leg kinematics. For robots that perform highly dynamic locomotion with long aerial phases, the filtering approaches using only proprioceptive sensors can result in incorrect or biased velocity estimation [6].
@@ III-A Moving Horizon Estimation
A general description of the dynamic system involves a nonlinear difference equation of its dynamics:
$x^{+}$ $=f(x,u)+w,$ (1)
and a measurement equation:
$y$ $=h(x)+v,$ (2)
@@
where $p(x_{[0,T]}|y_{[0,T]})$ is the posterior probability of the state trajectory $x_{[0,T]}$, and $(\cdot)^{*}$ denotes the optimal solution. When system is observable, the MAP problem (3) can be solved via the FIF formulation [17], with model descriptions in (1) and (2) being formulated as state constraints:
@@
${\hat{\mathcal{X}}_{[T-N,T]}}=$ $\underset{\mathcal{X}_{[T-N,T]}}{\text{argmin}}\ \Gamma(x_{T-N})+{\sum}_{k=T-N}^{T-1}{w_{k}^{\intercal}Q_{k}^{-1}w_{k}}\textstyle$
@@
As illustrated in Fig. 2, the arrival cost $\Gamma(x_{T-N})$ serves as an approximation metric to FIF, summarizing previous information prior to the current window. The arrival cost is calculated as described in (7), where the prior mean and covariance are typically calculated via EKF [16].
@@ III-B1 Sensors Models
$\tilde{a}$ $=a+b_{a}+\delta_{a},$ (9)
$\tilde{\omega}$ $=\omega+b_{\omega}+\delta_{\omega},$ (10)
@@
$b_{a}^{+}$ $=b_{a}+\delta_{b_{a}},$ (11)
$b_{\omega}^{+}$ $=b_{\omega}+\delta_{b_{\omega}},$ (12)
@@
$\tilde{\alpha}$ $=\alpha+\delta_{\alpha},$ (13)
$\dot{\tilde{\alpha}}$ $=\dot{\alpha}+\delta_{\dot{\alpha}}.$ (14)
@@ III-B2 Process Model
$\mathbf{x}^{+}=\begin{bmatrix}p+v\Delta t+\frac{1}{2}\big{(}R_{\text{WB}}(\tilde{a}-b_{a})+\text{g}\big{)}\Delta t^{2}\\ v+\big{(}R_{\text{WB}}(\tilde{a}-b_{a})+\text{g}\big{)}\Delta t\\ \zeta\big{(}(\tilde{\omega}-b_{\omega})\Delta t\big{)}\otimes q\\ p_{\text{foot}}\\ b_{a}\\ b_{\omega}\end{bmatrix}+\delta_{\mathbf{x}},$ (17)
where $\otimes$ represents the quaternion multiplication, $\zeta(\cdot)$ stands for the exponential map from the error rotation to the error quaternion:
@@
$\delta_{\mathbf{x}}\sim\mathcal{N}(0,Q_{\mathbf{x}})$ denotes the process noise; for detailed explanations, see [2].
@@ III-B3 Measurement Model
$y_{p}$ $=p-p_{\text{foot}}+\delta_{y_{p}},$ (19)
$\tilde{y}_{p}$ $=-R_{\text{WB}}fk(\tilde{\alpha}),$ (20)
@@ III-C Visual Odometry
Visual Odometry (VO) measures the pose of the robot in the camera frame $C$ by tracking features in the images from onboard cameras. Without loop closure [18], the VO output is interpreted as the incremental homogeneous transformation $\mathbf{T}_{\text{C}ij}$ between consecutive camera frames $C_{i}$ and $C_{j}$:
$\tilde{\mathbf{T}}_{\text{C}_{ij}}=\mathbf{T}_{\text{WC}_{i}}^{-1}\mathbf{T}_{\text{WC}_{j}}+\delta_{\text{vo}},$ (23)
where $\mathbf{T}_{\text{WC}_{i}}$ and $\mathbf{T}_{\text{WC}_{j}}$ are the homogeneous transformations from the world frame $W$ to the camera frame $C$ at time $i$ and $j$. The VO measurement is corrupted by Gaussian noise $\delta_{\text{vo}}\sim\mathcal{N}(0,Q_{\text{vo}})$. With known fixed transformation $\mathbf{T}_{\text{BC}}$ from the IMU/body frame $B$ to the camera frame $C$, the transformation of the body frame $\mathbf{T}_{\text{B}_{ij}}$ is represented as:
@@
In our decentralized estimation framework, (24) will be used to construct state constraints on incremental body displacement in the MHE in Section IV.B, and (25) corrects the orientation estimation in the EKF in Section IV.A.
@@ IV Decentralized estimation Framework
Considering the nonlinearity of the legged state dynamics and the extensive number of states and constraints involved, the corresponding nonlinear MHE poses significant computational challenges. We thus propose a novel approach to decentralize the nonlinear estimation of the floating base into a nonlinear orientation estimation using EKF and a linear velocity estimation using MHE, both of which then can be solved online. Similar decoupling has also been successfully applied for state estimation using KFs on the HRP2 humanoid [19], the Cassie biped [20], and MIT Cheetah 3 quadruped [21].
IV-A Orientation Estimation
@@ IV-A2 Measurement Model
$y_{a}$ $=R_{\text{WB}}^{T}\text{g}+\kappa\delta_{a},$ (27)
$\tilde{y}_{a}$ $=\tilde{a},$ (28)
@@
$y_{qc}$ $=q+\delta_{y_{qc}},$ (29)
$\tilde{y}_{qc}$ $=\text{rots}(\tilde{\mathbf{T}}_{\text{WB}}),$ (30)
@@
IV-B Position and Velocity Estimation
@@ IV-B3 Physical Constraint
The foot-ground contact is modeled as a deterministic constraint, i.e., the foot does not slip on the ground, instead of a stochastic model in EKF[2]. This model of deterministic state constraint will further demonstrate the capability of the MHE to handle constraints. When the foot is in stable contact with the ground,
@@
Note that this is a simplification of the general Linear Complementary Problem (LCP) formulation [25] of rigid body contact. It works well when the control enforces the foot contact to remain static during locomotion.
Figure 3: VO Synchronization and Interpolation.
@@ IV-B4 Synchronization and Visual Odometry Constraint
$y_{c_{k}}$ $=p_{k+1}-p_{k}+\delta_{c_{k}},$ (36)
$\tilde{y}_{c_{k}}$ $=\tilde{p}_{c_{k}},$ (37)
@@ IV-B5 Linear MHE
$\mathcal{X}_{[T-N,T]}$ $=\begin{bmatrix}\mathbf{x}_{[T-N,T]}\ \delta_{x_{[T-N,T-1]}}\ \delta_{c_{[T-N,T_{\text{vo}}-1]}}\ \delta_{y_{[T-N,T]}}\end{bmatrix}$
@@ V-B Arrival Cost Computation
We present the computation of arrival cost at time $T=N+1$; the consecutive computation follows the example with the window shifting forward. When $T\leq N$, the MHE is equivalent to FIF, and the arrival cost is equal to the prior cost as in (7), which can be written as $\Gamma(\mathcal{X}_{0})=\frac{1}{2}\mathcal{X}_{0}^{\intercal}M_{0}\mathcal{X}_{0}+m_{0}^{\intercal}\mathcal{X}_{0}$. At $T=N+1$, without marginalization, MHE with a window size of $N+2$ is equivalent to the FIF:
@@
where $J(\mathcal{X}_{[0,T]})$ is the quadratic cost of the (FIF) that can be written into the generalized form as in (41):
@@
The oldest optimization variable $\mathcal{X}_{0}=[\mathbf{x}_{0}\ \delta_{\mathbf{x}_{0}}\ \delta_{c_{0}}\ \delta_{y_{0}}]$ should be marginalized from (44) to maintain a fixed window size of the MHE (38), and the related constraints of $\mathcal{X}_{0}$ are incorporated to the arrival cost $\Gamma(\cdot)$. Based on the optimization structure of (44), it can be rewritten as:
@@
results in the ideal arrival cost $\Gamma(\mathcal{X}_{1})^{*}$ in the MHE (38) at $T=N+1$. This also eliminates the first row of constraints in (48), therefore, converting (47) to the MHE at $T=N+1$ as in (38). The closed-form solution of this arrival cost $\Gamma(\mathcal{X}_{1})$ is calculated using the KKT equations of the QP (44):
@@
where $\mathcal{L}=\begin{bmatrix}\mathcal{X}&\lambda\end{bmatrix}$; (51) is reordered from (43) w.r.t. the order of $\mathcal{L}_{[0,T]}$. Solving the Schur complement of the top-left corner of the KKT matrix $\mathbf{K}$ w.r.t. $\mathcal{L}_{0}$ yields the KKT equation of MHE in (38) that is marginalized from (44) [26]. Converting the resultant KKT equation back to the associated QP, that is, the MHE (38) at time $T=N+1$, we get the closed-form expression of its arrival cost $\Gamma(\mathcal{X}_{1})$:
@@
$m_{1}$ $=-k_{1}+K_{1,0}K_{0,0}^{-1}k_{0}.$ (54)
The consecutive marginalization process at a new time step follows the example at time $T=N+1$: the closed-form expression of $\Gamma(\mathcal{X}_{T-N})$ in (38) at time $T$ is obtained from the KKT condition of the pre-marginalized MHE of a window size of $N+2$ at time $T$. For the equality-constrained linear MHE, the above formulation entails no approximation, resulting in an MHE maintaining the optimality of FIF.
@@
Figure 5: Block Diagram of the Decentralized Estimation for PogoX.

### v3

@@ I Introduction
Due to the inherently underactuated nature of legged robots, an accurate and high-frequency estimation of the torso velocity and orientation is essential for the control of dynamic locomotion [2], [3]. This is a ubiquitous requirement for all dynamic legged robots regardless of how many legs they have. Extended Kalman filter (EKF) and its variants [2, 4, 5] are commonly used to solve this problem. These methods provide an accurate recursive solution for the fusion of onboard proprioceptive sensors including Inertial Measurement Units (IMU) and joint encoders. However, when working with locomotion behaviors associated with longer aerial phases, measurements from leg kinematics cannot be used to correct the floating base states. With only inertial sensors, the estimation of velocities can quickly diverge. Additionally, the yaw drift is inevitable when only using proprioceptive sensors [2]. To address this problem, several works [6, 7] have included exteroceptive sensors such as cameras and Lidar within the filtering methods. However, [6] primarily focused on pose drift correction, and [7] integrated vision and IMU into a single measurement, which compromised the robustness of the estimation due to potential correlations.
@@
To address this limitation, smoothing method that utilizes windowed optimization to include a history of sensor measurements is advantageous since multirate sensor data can be fused in the window. For instance, [3] utilizes a factor graph method, which tightly couples vision with IMU and encoder measurements through preintegration. However, in robotic implementations, factor graph-based methods typically have high computational demands, making their estimation rate insufficient for providing optimization results for real-time control. Separate EKFs are often employed in between the factor graph optimization to provide high-frequency estimates [3]. Another smoothing method, Moving Horizon Estimation (MHE), known as the dual problem of Model Predictive Control (MPC), provides an alternative formulation of the windowed optimization. Similar to MPC [8], when the dynamics is nonlinear, the MHE becomes a nonconvex optimization which yields demanding computation. Inspired by MHE, [9] proposed a Quadratic Program (QP) based state estimator using full-body dynamics with state constraints, to estimate contact force and joint torques. An inequality-constrained MHE was explicitly applied in [10] for a bipedal robot, demonstrating its ability to handle constraints and non-Gaussian noises. However, only proprioceptive sensors were considered in [10], and measurements prior to the smoothing window were directly discarded, which compromised the optimality of the estimation.
@@
Figure 1: The proposed estimation framework (bottom) and its application for realizing dynamic hopping on PogoX (top).
@@
In this paper, we aim to develop a MHE based framework that achieves multirate sensor fusion of both proprioceptive and exteroceptive sensors, providing fast and accurate estimation for legged robot control. The computation burden of nonlinear optimization of the canonical MHE is addressed by the decentralization of the orientation and velocity estimation: the nonlinear orientation estimation relies on an EKF, whereas the linear velocity estimation is solved via the MHE. The framework is illustrated in Fig. 1, and it is applicable to legged robots that have IMUs, leg joint encoders, and cameras. The work has the following contributions:
  • A novel state estimation framework, integrating a lightweight orientation EKF with a linear MHE for velocity estimation, to seamlessly fuse multirate proprioceptive and exteroceptive sensors while incorporating state constraints in real-time.
@@
  • Evaluations of the proposed decentralized estimators on various legged robotics systems, including the unipedal hopping robot PogoX [12], bipedal robot Cassie [4], and quadrupedal robot Unitree Go1 [13], with software implementation open-sourced at [14].
@@ II Related Work
State-of-the-art methods for legged robot state estimation are categorized into filtering and smoothing approaches. For filtering approaches, they typically perform fusions of proprioceptive sensors (IMUs, force/torque sensors, and joint encoders) with high frequency (100 - 2000 Hz) using Kalman Filters (KFs). [2] showed that using EKF, the robot pose, velocity, and IMU bias could be estimated on quadrupedal robots. An Invariant-EKF [4] was proposed to improve the convergence of the estimation. The filtering approaches have also been widely implemented on bipedal robots [5]. While providing high-frequency velocity estimation, the estimation accuracy is strongly affected by the availability of foot contact and accurate leg kinematics. For robots that perform highly dynamic locomotion with long aerial phases, the filtering approaches using only proprioceptive sensors can result in incorrect or biased velocity estimation [7].
@@ III-A Moving Horizon Estimation
A general description of the dynamic system involves the nonlinear dynamics and measurement equation:
$x^{+}=f(x,u)+w,\quad y=h(x)+v,$ (1)
@@
where $p(x_{[0,T]}|y_{[0,T]})$ is the posterior probability of the state trajectory $x_{[0,T]}$, and $(\cdot)^{*}$ denotes the optimal solution. When system is observable, the MAP problem (2) can be solved via the FIF formulation [18], with model descriptions in (1) being formulated as state constraints:
@@
${\hat{\mathcal{X}}_{[T-N,T]}}=$ $\underset{\mathcal{X}_{[T-N,T]}}{\text{argmin}}\ \Gamma(x_{T-N})+\textstyle{\sum}_{k=T-N}^{T-1}{w_{k}^{\intercal}Q_{k}^{-1}w_{k}}\textstyle$
@@
The basic concept of the MHE is illustrated in Fig. 2: the arrival cost $\Gamma(x_{T-N})$ serves as an approximation metric to FIF [17], which imposes previous information prior to the current window to the mean and covariance of $x_{T-N}$. The arrival cost is calculated as described in (6), where the prior mean and covariance are typically calculated via EKF [17].
@@ III-B1 Sensors Models
$\tilde{a}=a+b_{a}+\delta_{a},\quad\tilde{\omega}=\omega+b_{\omega}+\delta_{\omega},$ (8)
@@
$b_{a}^{+}=b_{a}+\delta_{b_{a}},\quad b_{\omega}^{+}=b_{\omega}+\delta_{b_{\omega}},$ (9)
@@
$\tilde{\alpha}=\alpha+\delta_{\alpha},\quad\dot{\tilde{\alpha}}=\dot{\alpha}+\delta_{\dot{\alpha}}.$ (10)
@@ III-B2 Process Model
$\mathbf{x}^{+}=\begin{bmatrix}p+v\Delta t+\frac{1}{2}\big{(}R_{\text{WB}}(\tilde{a}-b_{a})+\text{g}\big{)}\Delta t^{2}\\ v+\big{(}R_{\text{WB}}(\tilde{a}-b_{a})+\text{g}\big{)}\Delta t\\ \zeta\big{(}(\tilde{\omega}-b_{\omega})\Delta t\big{)}\otimes q\\ \begin{bmatrix}p_{\text{foot}}^{\intercal}&b_{a}^{\intercal}&b_{\omega}^{\intercal}\end{bmatrix}^{\intercal}\end{bmatrix}+\delta_{\mathbf{x}},$ (13)
where $\otimes$ denotes the quaternion multiplication, $\zeta(\cdot)$ is the map from the error rotation to the error quaternion:
@@
$\delta_{\mathbf{x}}\sim\mathcal{N}(0,Q_{\mathbf{x}})$ denotes the process noise.
@@ III-B3 Measurement Model
$y_{p}=p-p_{\text{foot}}+\delta_{y_{p}},\quad\tilde{y}_{p}=-R_{\text{WB}}fk(\tilde{\alpha}),$ (15)
@@ III-C Visual Odometry
Visual Odometry (VO) measures the robot pose in the camera frame $C$ by tracking features in the images from onboard cameras. Without loop closure [19], the VO output is interpreted as the incremental homogeneous transformation between consecutive camera frames $C_{i}$ and $C_{j}$: $\mathbf{T}_{\text{C}ij}=\mathbf{T}_{\text{WC}_{i}}^{-1}\mathbf{T}_{\text{WC}_{j}}+\delta_{\text{vo}}$, where $\mathbf{T}_{\text{WC}_{i}}$ and $\mathbf{T}_{\text{WC}_{j}}$ are the homogeneous transformations from the world frame $W$ to the camera frame $C$ at time $i$ and $j$. The VO measurement is corrupted by noise $\delta_{\text{vo}}\sim\mathcal{N}(0,Q_{\text{vo}})$. With known fixed transformation $\mathbf{T}_{\text{BC}}$ between IMU/body frame $B$ and camera frame $C$, the transformation of the body frame $\mathbf{T}_{\text{B}_{ij}}$ is represented as:
@@
In our implementation, the initial VO frame is aligned with the IMU frame when the VO is activated, assuming the robot starts in a static pose. In the proposed framework, (18) will be used to construct state constraints on incremental body displacement in the MHE in Section IV.B, and (19) corrects the orientation estimation in the EKF in Section IV.A.
@@ IV Decentralized estimation Framework
Considering the nonlinearity of the legged state dynamics and the extensive number of states and constraints involved, the corresponding nonlinear MHE poses significant computational challenges. We thus propose a novel approach to decentralize the nonlinear estimation of the floating base into a nonlinear orientation estimation using EKF and a linear velocity estimation using MHE, both of which can then be solved online. Similar decoupling has been successfully applied for estimation using KFs on the HRP2 humanoid [20], the Cassie biped [21], and MIT Cheetah 3 quadruped [22].
IV-A Nonlinear Orientation Estimation
@@ IV-A2 Measurement Model
$y_{a}=R_{\text{WB}}^{T}\text{g}+\kappa\delta_{a},\quad\tilde{y}_{a}=\tilde{a},$ (21)
@@
$y_{qc}=q+\delta_{y_{qc}},\quad\tilde{y}_{qc}=\text{rots}(\tilde{\mathbf{T}}_{\text{WB}}),$ (22)
@@
IV-B Linear Position and Velocity Estimation
@@ IV-B3 Physical Constraint
The foot-ground contact is modeled as a deterministic constraint, i.e., the foot does not slip on the ground. This deterministic state constraint model will further demonstrate the capability of the MHE to handle constraints. When the foot is in stable contact:
@@
Note that this is a simplification of the Linear Complementary Problem (LCP) formulation [26] of rigid body contact. It works well when the control enforces the static foot contact, and eliminates the need for extensive tuning of the sliding covariance. When walking on slippery terrain, the stochastic contact model in EKF [2] can still be enforced in the MHE.
Figure 3: The VO measurements are synchronized with the nearest IMU frames and interpolated at all IMU frames as the motion constraints.
@@ IV-B4 Synchronization and Visual Odometry Constraint
$y_{c_{k}}=p_{k+1}-p_{k}+\delta_{c_{k}},\quad\tilde{y}_{c_{k}}=\tilde{p}_{c_{k}},$ (28)
@@ IV-B5 Linear MHE
$\mathcal{X}_{[T-N,T]}$ $=\begin{bmatrix}\mathbf{x}_{[T-N,T]}\ \delta_{x_{[T-N,T-1]}}\ \delta_{c_{[T-N,T_{\text{vo}}-1]}}\ \delta_{y_{[T-N,T]}}\end{bmatrix},$
@@ V-B Arrival Cost Computation
We present the computation of arrival cost at time $T=N+1$; the consecutive computation follows the example with the window shifting forward. When $T\leq N$, the MHE is equivalent to FIF, and the arrival cost is equal to the prior cost (6), which can be written as $\Gamma(\mathcal{X}_{0})=\frac{1}{2}\mathcal{X}_{0}^{\intercal}M_{0}\mathcal{X}_{0}+m_{0}^{\intercal}\mathcal{X}_{0}$. At $T=N+1$, the pre-marginalized MHE, that is the “previous MHE” problem at $T-1$ with additional measurements at current time $T$, can be written as:
@@
which has a window size $N+2$; it is equivalent to the FIF. $J(\mathcal{X}_{[0,T]})$ is the quadratic cost of the pre-marginalized MHE that can be written into the generalized form as in (32):
@@
The oldest optimization variable $\mathcal{X}_{0}=[\mathbf{x}_{0}\ \delta_{\mathbf{x}_{0}}\ \delta_{c_{0}}\ \delta_{y_{0}}]$ should be marginalized from (35) to maintain a fixed window size of the MHE (29), and the related constraints of $\mathcal{X}_{0}$ are incorporated to the arrival cost $\Gamma(\mathcal{X}_{1})$. Based on the optimization structure of (35), it can be rewritten as:
@@
results in the ideal arrival cost $\Gamma(\mathcal{X}_{1})^{*}$ in the current MHE (29). This also eliminates the first row of constraints in (39), therefore, converting (38) to the current MHE (29). The closed-form solution of this arrival cost $\Gamma(\mathcal{X}_{1})$ is calculated using the KKT equations of the pre-marginalized MHE (35):
@@
where $\mathcal{L}=\begin{bmatrix}\mathcal{X}&\lambda\end{bmatrix}$; (42) is reordered from (34) w.r.t. the order of $\mathcal{L}_{[0,T]}$. Solving the Schur complement of the top-left corner of the KKT matrix $\mathbf{K}$ w.r.t. $\mathcal{L}_{0}$ yields the KKT equation of the current MHE (29) that is marginalized from the pre-marginalized MHE (35) [27]. Converting the resultant KKT equation back to the associated QP, that is, the current MHE in the form of (29), we get the closed-form expression of its arrival cost $\Gamma(\mathcal{X}_{1})$:
@@
$m_{1}$ $=-k_{1}+K_{1,0}K_{0,0}^{-1}k_{0},$ (45)
which does not require solving the KKT equation (42) or the pre-marginalized MHE (35). The consecutive marginalization process at a new time step follows the example at time $T=N+1$: the closed-form expression of $\Gamma(\mathcal{X}_{T-N})$ in (29) at time $T$ is calculated using (43), with (44) and (45) obtained from the KKT condition of the pre-marginalized MHE. For the equality-constrained linear MHE, the above formulation entails no approximation, resulting in an MHE maintaining the optimality of FIF. The marginalization is also computationally efficient as it only requires the inversion of $K_{0,0}$. With the arrival cost computed in closed form, any QP solver can be used to solve the MHE problem online.
@@
Figure 5: Block Diagram of the Decentralized Estimation on PogoX.